OpenStudy (anonymous):

we have two curve, f(x) and g(x), with f(a)=g(a), we want to make a transition from the f(x) to g(x) at (a,f(a))=(a,g(a)) , what kind of transition we can get if f'(a) not equal to g'(a)

OpenStudy (valpey):

This transition is continuous because f(a)=g(a) but it is not differentiable at a.

OpenStudy (anonymous):

i don't understand what this question ask. i know it's about expansion

OpenStudy (valpey):

Transitions can be discontinuous (when f(a)<>g(a)), they can be continuous but not differentiable (when f(a)=g(a) but f'(a)<>g'(a)) or they can be continuous and differentiable (when f(a)=g(a) and f'(a)=g'(a)). There might be cool words or names for these types of transitions.

OpenStudy (anonymous):

divergence and convergence?